# A Semi-Supervised Learning Method for the Identification of Bad Exposures in Large Imaging Surveys

### v1

@@ 2.5 Extinction Cut
Because the original Key Projects of the DESI experiment (DESI Collaboration et al., 2016), for which the LS was produced, focused on mainly extragalactic sources, the imaging catalog used for targeting also prioritizes sky areas with lower Galactic extinction to minimize the effect of dust attenuation. High-extinction exposures usually have dust emission features that will interfere with the training of our ML model, because the dust emission in the exposures generates additional (unwanted) features. Therefore, to minimize the influence of Galactic dust, we applied an extinction cut to exclude regions where $E(B-V)\geq 0.04$ (Schlegel et al., 1998), further down-selecting exposures for training.
@@ 2.6 Categories of Bad Exposures
  • non-optimal exposure: The exposure is underexposed or some similar exposure-related issue. This data is from the DECaLS list, where the keyword expfactor is included in a comment and is much less than 1.
  • Ghost/Scatter: Ghosting or scattered light is present in an image. We follow the convention of the DES bad exposure list to combine these two effects into one category. These effects are usually caused by very bright sources within or near a given field.
@@
  • Bad CCD: The CCD image has no data or has an amplifier issue or has a hot spot (i.e., pixel-level defects).
@@
  • Wonky: This category includes problems in exposures with patterns that do not belong to any other category and can be caused by miscellaneous reasons.
  • Telescope moving / tracking failure: The telescope tracking failed or the telescope unexpectedly moved.
  • Out of focus: The exposure was out of focus.
@@ 2.7 Balancing the Dataset
2.8 Final Training and Testing Samples
The final training and testing dataset contains two parts: the combined bad exposure list and the sample of good exposures. The final bad image list combines the DECam bad exposure list and the human-expert list from DES and DELVE. The total number of labeled bad exposures is 73,990, and detailed numbers from each source in each category are shown in Table 2.
Bad Category DECaLS list DES + DELVE from both Total
saturated 778 0 0 778
clouds or bad transparency 992 22331 15 23338
psf 2359 4781 501 7641
non-optimal exposure 3194 0 0 3194
Ghost/Scatter 0 15755 1 15756
NObjects 0 10093 0 10093
Bad CCD 0 2621 2 2623
Noise 0 2406 61 2467
Wonky 0 667 0 667
Telescope Moving 725 1403 601 2729
out of focus 2160 0 180 2340
Total 10208 60057 1361 73990
Table 2: The categories of the labeled bad exposures with the number of CCD images in each category from each dataset source. “DECaLS list” and “DES+DELVE” columns list data only from that source of exposures, exclusively. The “from both” column lists the exposures that are identified in both lists. As discussed in Section 2.1 and 2.2, not all categories are present in both lists.
@@ 3.1 Pipeline Overview
Our pipeline starts by feeding an image from the image dataset into the Deep Neural Network to generate an embedding with 768 dimensions (see Section 3.2). Some dimensions contain redundant or unstructured information (i.e., noise). We use post-processing to mitigate this noise to improve the overall prediction performance. The embedding post-processors include a scaler, a dimension reducer and a feature selector (Maharana et al., 2022). Output embeddings from the ViT are run through these post-processors before being passed to a kNN classifier (see Section 3.3). For the final step, the kNN classifier takes the post-processed embeddings and assigns a label with a probability based on the distance to the nearest neighbors in the training dataset. By comparing the probabilities and the labels, we can determine the qualities of the image from two perspectives: 1. whether the image is good or bad; 2. if the image is bad, which bad category it belongs to. Our end-to-end pipeline is illustrated in Figure 4.
[…]
@@ 3.4 Training and Hyperparameter Optimization
We employed the HPO technique (e.g., Franceschi et al., 2024) to search for the optimal hyperparameters for the post-processor and the kNN classifier. In particular, we defined ranges (for continuous variables) and options (for discrete variables) for hyperparameters and used the Random Search with Sucessive Halving (RSSH) algorithm (Jamieson & Talwalkar, 2016; Li et al., 2018) to search for the optimal hyperparameters. The ranges and options are chosen to be broad to let the RSSH algorithm explore for more globally optimal parameters. We also used 3-fold CV to evaluate the performance of the pipeline to ensure that our pipeline can be well generalized to a new dataset.
@@
where $p_{i}$ and $q_{i}$ are the $i$-th dimension of the two embeddings $p$ and $q$ summed over $dim$ dimensions. “Using 7 neighbors” in this sense means that the label and probability of each test image is determined by the distance to the 7 nearest neighbors in the training set.
@@ 3.5 Inference
After training and evaluation, we applied our pipeline to exposures that appear in DR11 of the LS (see Section 2) to infer which exposures might belong to the categories listed in Table 2. DR11 contains 283,633 exposures, so the inference step is somewhat computationally taxing. We therefore split the DR11 exposures into multiple batches and conducted the inference step in parallel on the National Energy Research Scientific Computing (NERSC) Center’s Perlmutter [^10] https://docs.nersc.gov/systems/perlmutter/architecture supercomputer using 2 nodes with 4 GPUs each. Fortunately, since the inference step doesn’t have to share gradients, it is embarrassingly parallelizable, and has very little communication overhead. Therefore, we can rapidly process many images in parallel with little-to-no performance degradation.
@@ 4.1 Clustering Analysis
We focus on using the training dataset to evaluate the clustering results, as it has known labels. The separation of the clusters is most easily visualized after running the embeddings through the dimension reduction process, the results of which are shown in Figure 5. Generally, well-separated and tightly-bound clusters indicate the model has learned enough features to distinguish different categories. We want to emphasize that plots such as Figure 5 – t-Distributed Stochastic Neighbor Embedding (t-SNE; van der Maaten & Hinton, 2008) plots are only useful for qualitative analysis, and that the displayed clustering results have been collapsed over many dimensions. This means that the axes of the figures, and the coordinates of each single data point, are not intuitively quantitatively interpretable in and of themselves.
@@
Figure 5: The clustering of the embeddings generated by the ViT model for the training dataset. The embeddings are processed through the same data processors described in Section 3.4, and they are further dimensionally reduced using the t-SNE method (van der Maaten & Hinton, 2008) to help with visualization. The grey dots represent all training datasets, the blue dots depict “good” exposures, and the orange dots highlight the labeled exposures for each “bad” category. The top-left panel displays the good exposures, and the other panels show the bad exposures. The axes of the figures correspond to two dimensions of the t-SNE-reduced embeddings, and are not physically interpretable.
@@
Figure 5 demonstrates that there is a clear clustering structure for the training dataset. There are several distinct areas that represent different categories of exposures. Most of the good exposures occupy the central parts of the space, while the bad exposures separate out across many different clusters. The good exposures typically contain various noticeable pattens that the model may pay attention to in a manner that a human visual inspector would not. Most prominently, emission from Galactic dust is easily identified by the model as a distinct feature. The model therefore incorrectly identifies dusty exposures as “bad”. We alleviate this issue for our final results by applying an extinction cut to the “good” exposure candidates to avoid dusty regions as described in Section 2.5.
@@
The wide spread of the “good” exposures indicates that the corresponding features are very diverse and span an extensive range of the embedding space, as one would expect from a large imaging survey. For the “bad” exposures, the “ghost” exposures and the exposures in the “NObjects” category are spread out and overlap with the clusters of “good” exposures. It is worth remembering that Figure 5 displays a two dimensional reduction of the embeddings, so overlapping clusters might actually be quite distinct in higher dimensions. We compare the “good” and “ghost” exposures in the next sub-section to better characterize the separation of the clusters in higher dimensions and to illustrate the capability of the model to distinguish different features of the images.
@@ 4.2 Classification Performance
Good 0.82 0.82 0.82 222
Saturated 0.98 0.97 0.98 122
Cloud_transparency 0.97 0.99 0.98 958
PSF 0.94 0.93 0.93 389
Nonoptimal_exp 0.94 0.96 0.95 204
Ghost_Scatter 0.92 0.80 0.85 222
NObjects 0.80 0.91 0.85 77
Bad_CCD 0.97 0.95 0.96 388
Noise 0.94 0.99 0.96 138
Wonky 0.71 1.00 0.83 12
Telescope_Moving 0.98 0.91 0.94 101
Out_of_focus 0.98 0.99 0.98 271
Table 3: The classification performance evaluated using three scores. The rightmost column (“Support”) records how many test images in a given category were used in the classification evaluation. This table is generated from the classification_report function in the scikit-learn package.
@@
The good performance demonstrated by our model across multiple evaluation metrics is due to several aspects of our approach. Fundamentally, we have confirmed in Section 4.1 that our model is able to recognize and distinguish different patterns. Furthermore, the 11 bad categories we use as labels are sufficiently representative to cover a wide range of photometric and technical issues. This may also imply that our model should be generalizable to other telescopes. Lastly, we applied a cut of $>0.9$ in probability for our classification metrics, meaning we only focus on the most problematic exposures, which will inevitably improve the accuracy of our technique.
@@ 5.1 Comparison between our pipeline and legacypipe for LS DR9 and DR10
[…]
Figure 10: Two exposures in DR11 that are identified by our pipeline. The reason given by our ML for the exposure on the left is Ghost_Scatter and for the right one is Telescope_Moving. The classification result is verified and confirmed by human experts.
Figure 11: Exposure 260867 from the DECaLS (in the $g$ band). The LS DR9 included this exposure, while our pipeline flagged this exposure as bad in the Ghost_Scatter category. Human experts later verified this exposure as bad during visual inspection. Several pupil ghosts are present in this exposure, which are likely caused by the bright star in the top-right corner.
Figure 12: Exposure 1302430 (in the $i$-band). This is an example of an image that was misclassified as bad by our pipeline. Our pipeline classified this image as being in the Telescope_Moving category. However, human experts later evaluated this exposure as good during visual inspection. The misclassification is likely due to the two parallel and proximate satellite trails, which are a novel feature that might mimic movement of the telescope.
@@ 6 Conclusion
We ran the pretrained ViT to generate embeddings that were later used to train the kNN classifier. The performance of our ViT+kNN pipeline was evaluated in two steps. We first evaluated the structure of the embeddings via clustering analysis to understand whether the model can successfully recognize and separate features that correspond to bad exposures. We then evaluated the classification performance of the kNN using several metrics, including accuracy and F1 score. Our model achieved a good and consistent performance with $>80\%$ classification precision for most categories of bad exposures.
@@ Acknowledgements
Appendix A Data Availability Statement
@@
Appendix B Reproducibility Considerations of ML models

### v2

@@ 2.5 Extinction Cut
Because the original Key Projects of the DESI experiment (DESI Collaboration et al., 2016), for which the LS was produced, focused on mainly extragalactic sources, the imaging catalog used for targeting also prioritizes sky areas with lower Galactic extinction to minimize the effect of dust attenuation. High-extinction exposures usually have dust features that will produce patterns that interfere with the training of our ML model. Therefore, to minimize the influence of Galactic dust, we applied an extinction cut to exclude regions where $E(B-V)\geq 0.04$ (Schlegel et al., 1998), further down-selecting exposures for training.
@@ 2.6 Categories of Bad Exposures
  4. 4.
    Nonoptimal_exp: The exposure is underexposed or some similar exposure-related issue. This data is from the DECaLS list, where the keyword expfactor is included in a comment and is much less than 1.
  5. 5.
    Ghost_scatter: Ghosting or scattered light is present in an image. We follow the convention of the DES bad exposure list to combine these two effects into one category. These effects are usually caused by very bright sources within or near a given field.
  6. 6.
@@
  7. 7.
    Bad_CCD: The CCD image has no data or has an amplifier issue or has a hot spot (i.e., pixel-level defects).
  8. 8.
@@
  9. 9.
    Wonky: This category includes artifacts in exposures with patterns that do not belong to any other category and can be caused by miscellaneous reasons.
  10. 10.
    Telescope_moving: The telescope tracking failed or the telescope unexpectedly moved.
  11. 11.
    Out_of_focus: The exposure was out of focus.
[…]
figures/exp-347737-Saturated.png0.3(a) Saturated
@@
As discussed in Sec. 1, some issues, such as PSF, Out of focus, and Saturated can be detected by existing automated methods. We decided to retain these categories for two reasons. First, these well-characterized issues can be used to validate our pipeline. Second, because the class labels our model considers are exclusive, including well-characterized issues allows the model to utilize complementary information about less-well-characterized patterns, which can improve the performance of the model across all categories.
@@ 2.7 Balancing the Dataset
2.8 Summary
Our dataset draws from DECaLS imaging in LS DR9 and DR10, supplemented with exposures from DELVE. Problematic exposures are identified from four main sources, which we summarize below (see also Table 1):
  1. 1.
    The DECaLS Bad Exposure List, which we use to flag poor quality exposures, bad observing conditions, or instrument issues.
  2. 2.
    DES Expert Labels. These are visual inspections from the DES Y6 release.
  3. 3.
    DELVE Expert Labels. These are visual inspections from DELVE DR3.
  4. 4.
    LS pipeline flags. These are automated quality flags applied during LS pipeline processing, which we use to prune further potentially bad exposures from the good exposure candidate set.
We unified the labels from the DECaLS and human expert lists into the 11 categories described in Sec. 2.6, focusing on the most common issues that can be generalized across exposures. We reduced the dataset to these categories to improve model generalization, downsampled overrepresented classes, and balanced the good/bad exposure ratio for classification. The final labeled set contains 73,990 bad exposures and 10,000 good exposures.
@@ 3.1 Pipeline Overview
Our pipeline starts by feeding an image from the image dataset into the Deep Neural Network to generate an embedding with 768 dimensions. Some dimensions contain redundant or unstructured information (i.e., noise). We use post-processing to mitigate this noise to improve the overall prediction performance. The embedding post-processors include a scaler, a dimension reducer and a feature selector (Maharana et al., 2022). Output embeddings from the ViT are run through these post-processors before being passed to a kNN classifier. For the final step, the kNN classifier takes the post-processed embeddings and assigns a label with a probability based on the distance to the nearest neighbors in the training dataset. By comparing the probabilities and the labels, we can determine the qualities of the image from two perspectives: 1. whether the image is good or bad; 2. if the image is bad, which bad category it belongs to. Our end-to-end pipeline is illustrated in Figure 4.
ImageDatasetViTModelPost-processorClusteringAnalysiskNNClassifierlabel $\neq$ 0andProb >> 0.9?BadGoodBad CategoryProbabilitylabelYesNoAssign labelwith probability
@@ 3.2 Training and Hyperparameter Optimization
We employed the HPO technique (e.g., Franceschi et al., 2024) to search for the optimal hyperparameters for the post-processor and the kNN classifier. In particular, we defined ranges (for continuous variables) and options (for discrete variables) for hyperparameters and used the Random Search with Sucessive Halving (RSSH) algorithm (Jamieson and Talwalkar, 2016; Li et al., 2018) to search for the optimal hyperparameters. The ranges and options are chosen to be broad to let the RSSH algorithm explore for more globally optimal parameters. We also used 3-fold CV to evaluate the performance of the pipeline to ensure that our pipeline can be well generalized to a new dataset. Since our task is a multi-class classification problem, we used the overall accuracy (see Equation 1) as the metric to evaluate the performance of the pipeline during HPO – the higher the accuracy, the better the hyperparameter combination. The hyperparameter search space is summarized in Table 2.
Component Hyperparameter Search Space
Scaler StandardScaler, MinMaxScaler, Normalizer,
MaxAbsScaler, RobustScaler, PowerTransformer
PCA number of components 5, 10, 15, 25, 30, 50, 70, 100
VarianceThreshold 0, 0.001, 0.01
kNN number of neighbors 1, 3, 5, 7, 10, 20, 30
kNN distance metric (p) 1, 2, 3, 5
kNN leaf size 1, 5, 10, 15, 30, 35
Table 2: The hyperparameter search space for HPO. The bold values indicate the determined optimal hyperparameters.
@@
where $p_{i}$ and $q_{i}$ are the $i$-th dimension of the two embedding vectors $p$ and $q$ summed over dimdim dimensions (i.e. essentially the sum of the absolute differences across the vector components). “Using 7 neighbors” in this sense means that the label and probability of each test image is determined by the distance to the 7 nearest neighbors in the training set.
@@ 3.3 Inference
After training and evaluation, we applied our pipeline to exposures that appear in DR11 of the LS (see Section 2) to infer which exposures might belong to the categories listed in Table 4. DR11 contains 283,633 exposures, so the inference step is somewhat computationally taxing. We therefore split the DR11 exposures into multiple batches and conducted the inference step in parallel on the National Energy Research Scientific Computing (NERSC) Center’s Perlmutter [^10] https://docs.nersc.gov/systems/perlmutter/architecture supercomputer using 2 nodes with 4 GPUs each. Fortunately, since the inference step doesn’t have to share gradients, it can be run on completely independent GPUs. Therefore, we can rapidly process many images in parallel with little-to-no performance degradation.
@@ 4.1 Clustering Analysis
We focus on using the training dataset to evaluate the clustering results, as it has known labels. The separation of the clusters is most easily visualized after running the embeddings through the dimension reduction process, the results of which are shown in Figure 5. Generally, well-separated and tightly-bound clusters indicate the model has learned enough features to distinguish different categories. We want to emphasize that plots such as Figure 5 – t-Distributed Stochastic Neighbor Embedding (t-SNE; van der Maaten and Hinton, 2008) plots are only useful for qualitative analysis, and that the displayed clustering results have been collapsed over many dimensions. This means that the axes of the figures, and the coordinates of each single data point, are not intuitively quantitatively interpretable in and of themselves.
@@
Figure 5: The clustering of the embeddings generated by the ViT model for the training dataset. The embeddings are processed through the same data processors described in Section 3.2, and they are further dimensionally reduced using the t-SNE method (van der Maaten and Hinton, 2008) to help with visualization. The grey dots represent all training datasets, the blue dots depict “good” exposures, and the orange dots highlight the labeled exposures for each “bad” category. The top-left panel displays the good exposures, and the other panels show the bad exposures. The axes of the figures correspond to two dimensions of the t-SNE-reduced embeddings, and are not physically interpretable.
@@
Figure 5 demonstrates that there is a clear clustering structure for the training dataset. As a more quantitative metric, the “Silhouette Score” [^11] See https://scikit-learn.org/stable/modules/generated/sklearn.metrics.silhouette_score.html of the clustering is 0.23, showing some indications of clustering. There are several distinct areas that represent different categories of exposures. Most of the good exposures occupy the central parts of the space, while the bad exposures separate out across many different clusters. The good exposures typically contain various noticeable pattens that the model may pay attention to in a manner that a human visual inspector would not. Most prominently, extinction caused by Galactic dust is easily identified by the model as a distinct feature. The model therefore incorrectly identifies dusty exposures as “bad”. We alleviate this issue for our final results by applying an extinction cut to the “good” exposure candidates to avoid dusty regions as described in Section 2.5.
@@
The wide spread of the “good” exposures indicates that the corresponding features are very diverse and span an extensive range of the embedding space, as one would expect from a large imaging survey. Due to the high complexity of the images, the weak separation of clusters is mainly caused by the large scattering of “good” exposures. We will discuss how a probability cut in Sec. 4.2 can reduce the effect of this scattering. For the “bad” exposures, the “ghost” exposures and the exposures in the “NObjects” category are spread out and overlap with the clusters of “good” exposures. It is worth remembering that Figure 5 displays a two dimensional reduction of the embeddings, so overlapping clusters might actually be quite distinct in higher dimensions. We compare the “good” and “ghost” exposures in the next sub-section to better characterize the separation of the clusters in higher dimensions and to illustrate the capability of the model to distinguish different features of the images.
@@ 4.2 Classification Performance
Good 0.970 0.764 0.854 2411
Saturated 0.990 1.000 0.990 411
Cloud_transparency 0.977 0.990 0.981 3076
PSF 0.872 0.960 0.915 940
Nonoptimal_exp 0.925 0.970 0.947 599
Ghost_scatter 0.833 0.910 0.868 693
NObjects 0.824 0.960 0.886 283
Bad_CCD 0.949 0.980 0.963 1149
Noise 0.842 1.000 0.913 461
Wonky 0.814 1.000 0.899 49
Telescope_moving 0.804 0.950 0.870 284
Out_of_focus 0.974 0.990 0.981 874
Table 3: The classification performance evaluated using three scores. The rightmost column (“Support”) records how many test images in a given category were used in the classification evaluation. This table is generated from the classification_report function in the scikit-learn package. We re-ran our analysis with 10 sets of random seeds and found that the standard deviation of each metric is less than 0.025, indicating the stability of our model.
@@
The good performance demonstrated by our model across multiple evaluation metrics is due to several aspects of our approach. Fundamentally, we have confirmed in Section 4.1 that our model is able to recognize and distinguish different patterns. Furthermore, the 11 bad categories we use as labels are sufficiently representative to cover a wide range of photometric and technical issues. This may also imply that our model should be generalizable to other telescopes. Lastly, we applied a cut of >0.9>0.9 in probability for our classification metrics, meaning we only focus on the most problematic exposures, which will inevitably improve the accuracy of our technique.
@@ 6 Conclusion
We ran the pretrained ViT to generate embeddings that were later used to train the kNN classifier. The performance of our ViT+kNN pipeline was evaluated in two steps. We first evaluated the structure of the embeddings via clustering analysis to understand whether the model can successfully recognize and separate features that correspond to bad exposures. We then evaluated the classification performance of the kNN using several metrics, including accuracy and F1 score. Our model achieved a good and consistent performance with $80\%$ classification precision for most categories of bad exposures.
@@ Acknowledgements
Appendix A Final Training and Testing Samples
Our final training and testing dataset comprises two parts: the combined bad exposure list and the sample of good exposures. The final bad image list incorporates the DECam bad exposure list and the human-expert lists from DES and DELVE. The total number of labeled bad exposures is 73,990, and detailed numbers from each source in each category are shown in Table 4.
Bad Category DECaLS list DES + DELVE from both Total
saturated 778 0 0 778
clouds or bad transparency 992 22331 15 23338
psf 2359 4781 501 7641
non-optimal exposure 3194 0 0 3194
Ghost/Scatter 0 15755 1 15756
NObjects 0 10093 0 10093
Bad CCD 0 2621 2 2623
Noise 0 2406 61 2467
Wonky 0 667 0 667
Telescope Moving 725 1403 601 2729
out of focus 2160 0 180 2340
Total 10208 60057 1361 71626
Table 4: The categories of the labeled bad exposures with the number of CCD images in each category from each source. The “DECaLS list” and “DES+DELVE” columns include data exclusively from those sources. The “from both” column counts the exposures that are identified in both lists. As discussed in Sec. 2.1 and 2.2, not all categories appear in both samples. Images counted in this table are selected to avoid dust extinction, as discussed in Sec. 2.5.
Appendix B Sample Exposures Identified by the Model
[…]
Figure 10: Two exposures in DR11 that are classified as “bad” by our pipeline. Our ML labels the exposure on the left (right) as Ghost_Scatter (Telescope_Moving). These classifications have subsequently been verified by human experts.
Figure 11: Exposure 260867 from the DECaLS ($g$-band). The LS DR9 included this exposure, while our pipeline flagged it as bad in the Ghost_Scatter category. Human experts later confirmed that this exposure was bad during visual inspection. Several pupil ghosts are present in this exposure, which are likely caused by the bright star in the top-right.
Figure 12: Exposure 1302430 (ii-band). This is an example of an image that was misclassified as “bad” (in the Telescope_Moving category) by our pipeline. Human experts later evaluated this exposure as good during visual inspection. The misclassification is likely due to the two parallel and proximate satellite trails, which are a novel feature that might mimic movement of the telescope.
Appendix C Data Availability Statement
@@
Appendix D Reproducibility Considerations of ML models